1/27
Looks like no tags are added yet.
Random
Generates outcomes purely by chance
Probability
A number between 0 and 1 that describes the proportion of times the number would occur in a very long series of trials
Not a simulation
Law of Large numbers
Says if we observe more and more repetitions of any chance process, the proportion of times that a specific outcome occurs approaches a single value that is the true probablity
Simulation
The imitation of chance behavior based on a model that accurately reflects the situation
Preforming a simulation
1) Describe how to set up and use a random process to preform one trial of the simulation
2) Preform many trials (20 is good to get a estimate)
3) Use the results of your simulation to answer the question of interest
Preforming a simulation to estimate the probability of getting at least one run of 3 heads in a row in 10 flips of a fair coin→ correct simulation description
Assign the numbers 0 through 4 to “heads” and 5 through 9 to “tails” Read 10 one-digit numbers from a random digits table and count this simulation as a “success” if there is at least one run of 3 or more heads, repeat 500 times and divide the total number of “successes” by 500
A or B or AUB
Union of A and B→ will consist of all outcomes in events A, B or both
Event
The specific outcome or set of outcomes of a random phenomenon
ex: flip coin and #, roll dice and 2
Probability model
Lists all the possible outcomes of an events and their probabilities
Sample space (S)
The list of all possible outcomes
A and B
intersection of A and B→ consists of all outcomes common to both events
To find probability that an event occurs
P(A)= # of outcomes in event A/ total # of outcomes in event space
P(S)=1
The sum of probability for the sample space
Complement formula
P(A^ complement)=1-P(A)
The probability A does NOT occur
ex: P (at least 1)=1-P(None)
General addition rule
P(A or B)= P(AUB)= P(A)+P(B)-P(A and B)
Mutually exclusive (Disjoint events)
Events that do not overlap
The events have no outcomes in common and therefore can never occur together
If A and B are mutually exclusive then: P(A or B)= P(A)+P(B)
Sample w/ replacement
Card is returned to deck→ shuffled→ prob is the same
Sampling w/o replacement
Card is removed, prob changes
Union
The set of all outcomes that are A or B or both (Or) (P AUB)
Complement
A^C= not A→mutually exclusive/disjoint
Intersection
The set of all outcomes that are both A and B at the same time (and)
General addition rule
For any 2 events A and B
P(A or B)= P(A)+ P(B)-P(A and B)
Valid Prob. Models
-Prob of event is # between 0 and 1
-All possible outcomes together must have a probabilities that add up to one
-Probability that an event does not occur is 1-minus the probability that the event does occur
Question Wording
And=overlapping box
Given=use the total of whatever comes after your line as your denominator
Or=add the relevant categories (subtract the over lap if needed
P(hit)→total hits/ grand total
P(Post)→total post-season/grand total
P(hit and post)→ overlapping box/grand total
P(hit given post)→overlapping box/post season row total
conditional probability
The probability of event A under the condition that we know event B has occured
Prob of A given B→ P(A|B) → P (A and B)/P(B)
Prob of B given A → P(B|A)=P (A and B)/P(A)
Conditional probability and independent
When knowledge that one event has happened does not change the likelihood that another event will happen
P(A|B)=P(A) and P(B|A)=P(B)
General Multiplication rule
For any 2 events A and B (can be disjoint or not, dont overlap)
P(A and B)=P(A)xP(B|A)
For both of the two events to occur, first one must occur and then given that the first one has occurred, the second must occur
Tree diagram
All possibilities of what your doing
The probability of reaching the end of any complete branch is the product if the probabilities written in its segments
Multiply to the end